OpenStudy (anonymous):

What mass, in grams, of chlorine gas (Cl2) is contained in a 15.5 liter tank at 25.5 degrees Celsius and 1.98 atmospheres? Show all of the work used to solve this problem.

OpenStudy (anonymous):

So far I have PV=nRT (1.98)(15.5)= n(0.0821)(25.5) 30.69=n(2.09355) 14.65931 = n moles=grams/molecular mass grams = moles x mol mass g = 14.659 x 35.453 g =

OpenStudy (anonymous):

I just want to make sure it's right

OpenStudy (accessdenied):

Your temperature was given in degrees Celsius. But the Ideal Gas Law requires the absolute temperature in Kelvin, correct?

OpenStudy (accessdenied):

Also, the molecular mass of a single chlorine atom is 35.453 g/mol, but the question specifically uses diatomic chlorine gas (Cl_2). You have two chlorine atoms, so you should use twice that molar mass for a unit of chlorine gas. Other than that, the remaining steps look great. Just some little things that are easy to overlook!